OpenStudy (anonymous):

The length of a rectangle is 4 inches more than its width, and its perimeter is 36 inches. Find the width of the rectangle. 11?

OpenStudy (anonymous):

if w is the width,what is the length?

OpenStudy (anonymous):

\[P = 2(l + w)\]\[P = 2(w + 4 + w)\]\[36 = 2(2w + 4)\]\[\frac{36}{2} = \frac 22(2w + 4)\]\[18 = 2w + 4\]Solve for w and there is your width.

OpenStudy (anonymous):

7

OpenStudy (anonymous):

Yup :) Good job!